Change page style: 

GeMS News

You are here

December. 2016

GeMS laser update


In October 2015, a request for laser quotes was issued. A review led in February 2016 by a committee of Gemini laser project members and external specialists of lasers for Adaptive Optics proposed the selection of the quote presented by the Toptica Photonics AG. In March 2016, the NSF approved the procurement of a Toptica SodiumStar 20/2 Laser to equip the Gemini South Observatory. The laser successfully passed its Factory Acceptance Tests and is currently on its way to the Gemini South Observatory. We are aiming to install the new laser over the next semester and to switch to the Toptica laser mid-2017.

June. 2015

GeMS laser


During this semester, a team of specialists from Gemini North and South explored the options opened by new laser technologies to recover enough laser photon return to sustain GeMS long-term performances. Under the frame of a "laser feasibility study" project, this dedicated team found several candidates could indeed replace the current laser for an equivalent photon return. Their conclusions were recently approved by both STAC and governance Boards. Gemini will initiate a laser procurement next semester.

June. 2015

Semester 2015A runs


We managed with great effort to stabilize the laser power at an average of 31W during this full semester. Thanks to the laser team for their continuous effort! The March run including some guaranteed time for the ANU GSAOI team and was entirely run under clear sky and good seeing. GeMS sustained good stability and high performances leading to a very productive run overall. The May-June run, including some extended time for GSAOI Long Program 2, was more difficult due to poor atmospheric conditions. GeMS performed very well during the few episodes where the sky was clear and the seeing decent. However the rest of the run was plagued with a mix of clouds (5 nights entirely lost and 2 partially) and very poor seeing. The AO team managed to deliver some decent data during some of the poor seeing episodes.

Dec. 2014

December run 2014


The November run was cancelled due to laser issues. With the help of an external consultant, our engineers were able to achieve a stable 35W level. Classical nights with observers coming from ANU (Australia) and LAM (France) were very succesful. We made some good progress on the regular queue as well.

June-Sept. 2014

Winter Shutdown 2014


GeMS/GSAOI was removed from the telescope and brought back to the instrument lab during the winter shutdown. A comprehensive campaign of optical tests pointed to field stops misalignments in the LGSWFS. Re-alignment of the LGSWFS was consequently undertaken. Despite the bad weather, it was successfully tested during our September run.

June-Sept. 2013

Winter Shutdown


GeMS/GSAOI was removed from the telescope and brought back to the instrument lab for the winter shutdown. Work on GeMS covered a re-alignment of the NGS and LGSWFS, the BTO and the Laser. GSAOI was upgraded with new motors for the filter wheels. All systems were back on-sky in early September.

January to June 2013

System Verification and first queue semester


In the period from January 2013 to June 2013, GeMS was used around 8 night / month to run science. First the System Verification programs, and then the 13A programs were executed. A total of around 150 hours of science images has already been delivered by the system. SV data are publicly available at and a press release summarize the nice images already obtained:

26th of December to 1st of January 2013

Last GeMS run of 2012 - Starting some SV


During the very last run of 2012, we were able to start some of the SV observations. Among others, a very nice field was observed: the Orion bullets. Check the following press release for all the details:

November 2012 to December 2012

GeMS back on sky: objective ends of commissioning


Between November and December 2012, GeMS was working on sky for four commissioning runs. These runs were the last ones to finish commissioning, and start the transitions toward System Verification and operations for 13A. First runs will be dedicated to the Laser Guide Star Facility, while the next runs will be dedicated to performance optimizations, stabilization, and GSAOI science commissioning.

1st June 2012 - 1st November 2012

GeMS winter shutdown


For the second year, GeMS was in winter shutdown for the period June to November. This shutdown aimed at fixing and upgrading sub-systems, based on experience gathered on-sky. On the menu for this year: automation of motors on the laser bench, re-alignment of the laser bench, changing motors on the BTO, commissioning of new beam splitters in Canopus, full re-alignment of the Laser Guide Star WFS, potential upgrade for the Natural Guide Star WFS, and transmission measurement and cleaning of all optics.

20 Mar. 2012

Call for community commissioning targets


See the following call for a first opportunity to use GeMS/GSAOI.

See: For all the details

8-15 Mar. 2012

10th commissioning run: GeMS with GMOS-S


Due to issues with the cooling of GSAOI, we used GeMS with GMOS. Although the combination of GMOS and GeMS is not intended to be offered as a standard mode in the immediate future, this image is useful for commissioning/verification purposes and demonstrates the capabilities of GeMS over a broad spectral range. It also hints at what is possible with a ground-layer adaptive optics system. Good performance were obtained, with Full Width Half Max (FWHM) of as good as 0.08 arcsec in i-band.

12-20 Dic. 2011

7th commissioning run: GeMS at work.


This was (so far) the best run we had. The main goal of this run was to demonstrate that GeMS is working, and it is working well.

See: For all the details

9-15 Nov. 2011

6th commissioning run.


Back on sky for this 6th commissioning run. Goal is to check that everything that was working is still working, and that the new upgrades from the shutdown are operational.

Results: it was a very successful run, we achieved almost all the goals that we have set.
Among others:
- The new FSA works perfectly
- The LLT focus remote control as well
- Laser performance were stable
- Laser operations (spotters, plane, LCH) were smooth
- Canopus was super stable, and performed well, the new 2DMs config is fully integrated now
- Many software bugs, hardware issues, electronic glitch have been solved very efficiently on the fly
- The transition from shutdown to commissioning was remarkably smooth
That's promising for the next runs.


7&8 Nov. 2011

Canopus & GSAOI are respectively back at the telescope


After a five-month marathon or hard engineering work on both instruments, they are back at the telescope on schedule.
Below is a non-exhaustive list of upgrades that has been performed during that period:

The Fast-Steering Array (FSA) has been fully re-designed, re-constructed and re-implemented
The Laser Launch Telescope (LLT) has been equipped with a motorized focus screw
The Laser has been fully cleaned, re-aligned and re-optimized
All the mechanisms of Canopus have been re-optimized
Entrance window of GSAOI has been cleaned
The Beam Transfer Optics (BTO) has been at the heart of a large engineering effort, combining mechanical, electronic and software upgrades.
The Natural Guide Star (NGS) Wave-Front Sensor (WFS) of Canopus has been upgraded for a better throughput
And many more

In summary it has been a very intense period during which many activities were orchestrated in parallel, and which eventually result in a more reliable and better performing system.


1st Jun. 2011

Planned five-month rework period


In early June 2011, GeMS entered a planned five-month rework period and the 10th of June, Canopus was removed from the telescope and installed in the instrument lab of Cerro Pachon. This shutdown was planned to perform upgrades of subpackages that were deliberately put aside when we decided to go on the telescope end of last year. It also appeared to be the perfect opportunity to fix, repair and upgrade many systems based on the experience that was acquired on-sky, taking advantage of the winter time in Chile less favorable to AO observations.

GeMS is scheduled to be back at the telescope on November.


16-22 May. 2011

5th and last commissioning


5th and last commissioning of this first period. Unfortunatly it was a difficult commissioning run, extract from the GeMS blog:

Autopsy of a run - on May 24, 2011 2:39 PM
This past run was a nightmare. On top of the bad weather which was expected, we had a surprising number of major issues.
Laser / BTO / GSAOI cryocooler: no of those was cooperative to work at night. Adding to that the bad weather: clouds (2-3 nights of cirrus), high wind (1 night at about 15m/s, one night at 25-30m/s), bad seeing (we saw up to almost 3 arcsec).
The big disappointment was to not be able to do compensated images with GSAOI at all, notwithstanding the nice images we were aiming for.


15-21 April. 2011

4th commissioning run of 2011


4th commissioning run of 2011: when everything starts to work together and the first image came out of GeMS !


Here is the status of the system during this recording:
* LGS loop closed
* NGS loop closed with 3 TTGS
* Offloads: M1, M2, FSA, dithering

What is that exactly and what can I see?
* This is the MYST client Real-Time Display
* It fetches data packets from the RTC, displays raw and processed data at up to 30 frames/seconds
* The upper left image is dedicated to WFSs. At the beginning of the video, it shows the WFS pixels. The cross-shaped dominant feature is the signature of the Rayleigh (or L-shaped in the outer WFS). Later, some WFS statistics are displayed.
* The right hand side display the shape of the 3 deformable mirrors and Tip-Tilt mirror position (green spot), plus some statistics. Note how one can see the turbulence passing by, especially on DM0 (bottom image in right hand panel).
* The bottom left panel display the 3 TT WFS APD signal. Obviously, the TTWFS#3 looks at a much brighter source than the others.The signal in the 4 quadcell of each WFS may not be exactly equal because we correct only the average TT and not the Tilt anisoplanatism (yet to come). After a few seconds, this panel displays the Zernike coefficients rms and an estimation of r0.

More info here
And the december 2011 issue of the Gemini Focus.




21-27 March. 2011

Starting the 3rd commissioning run


Starting the 3rd commissioning run: goals are to finish all the LGSF functionalities and start the Canopus commissioning.
Succesfull run, a good summarize would be this extract from the GS NightLog: 2011 mar24-mar25

"Very good progress tonight. We went throught a lot of offloads debugging, and closing loops.
We can now close the FSA offload to keep the laser in front the LGSWFS.
We closed the LGS (with the 3 DMs and 5 WFSs) with decent gains.
We acquired 1 TipTilt star and closed the Tip-Tilt loop.
We worked out offload from the TTM to M2. Eventually we closed both LGS and NGS loops at the same time, with FSA and M2 offloads enabled."

For more info you can download a special Gemini Focus (March 2011)


20-26 Feb. 2011

Starting the 2nd run of the Laser Guide Star facility (LGSF) / GeMS commissioning.


Lot of progress has been done in the BTO control, constellation alignement and more propagation at Zenith and outside Zenith.
Highlight of this run is the first close loop with Canopus !

Extract from Canopus blog:
By Francois Rigaut on February 25, 2011 6:56 AM

We did it!
Tonight around 5:47AM, we closed the loop with the LGS WFS on the LGSs. Low gain, fat spots, no real turbulence compensation, but at least it was stable (we were closed for 10mn+) and correcting the static aberrations. We had enough photons at 200Hz. This is all good news, and looks good for the future commissioning runs.


22 Jan. 2011

First laser propagation


"In the early morning of January 22, 2011 at 4:38am, Chile Summer Time, a new era in high-resolution astronomy began with the successful propagation of a 5-star sodium laser guide star "constellation" in the skies over Cerro Pachon in Chile."

More details can be found at the following link:


10 Jan. 2011

Canopus installation on the Telescope


On Monday 10th of January, Canopus was installed on telescope. On the 11th, power was turned on. Canopus will now be characterized in this new environment for around 2months, before starting the night time commissioning


25 Nov. 2010

Laser high power spec achieves in LSE.


Vincent Fesquet and LMCT Laser engineer Zach Prezkuta have just spent 3 weeks of nearly non stop work in the laser Service Enclosure to bring the GS Laser back to its preshipment performance level. They succeeded pretty nicely, achieving between 50 to 60W most days during the last week and close to 65W on the day when Zach left back to Colorado. Wavelength locking was achieved a few days earlier. Other performance parameters such as beam quality and stability will be assessed over the next couple of weeks.


19 Nov. 2010

Canopus Transport to Cerro Pachon and installation in the instrument lab.


Canopus was transported from the lab in la Serena base facility to the instrument lab at the telescope in cerro Pachon. The preparation, packing and transportation took 3 days of work for a team of more than 10 engineers and technician. The transport went smoothly, and Canopus is now installed in the instrument lab, where it will be tested before its installation on the Telescope planned for the 10th of January.


14 Jul. 2010

GS laser installation on telescope


The GS laser passed its last Software reliability acceptance test last week and was officially accepted on Friday July 9, 2010. Preparations had been made so that the laser could be installed on the telescope immediately thereafter. Prep work at the summit occurred on Mon-Tue July 12-13 for the Laser Bench Enclosure (LBE) and Laser Electronic Enclosure (LEE) to be transported from the Pachon Instrument Lab clean room to the telescope and installed in their final home: the GS Laser Service Enclosure (LSE). The move took place today, on July 14, 2010, under Mike Sheehan's supervision and with help from a large section of the CP day crew. The move went pretty smoothly. Both the LBE and LEE were installed in the LSE by ~7pm tonight. Look at this podcast for more details.


24 Mar, 2010

Gemini South Laser delivery to Cerro Pachon


The GS laser arrived at 1pm today on the summit. All crates were offloaded by 2:30pm, and stored on the first floor where we will pick them up again on Monday for installation in the instrument lab. Everything went by the book.


Gemini Observatory Participants